What is the perimeter of rectangle WXYZ with coordinates W(3, -4), X(3, -6), Y(-1, -6), and Z(-1, -4)?
Answer by ikleyn(52786)

.

WX is vertical of the length of 2.

XY is horizontal of the length of 4.


The perimeter is 2*(2+4) = 2*6 = 12.     ANSWER

Solved, answered and explained.